OpenStudy (anonymous):

solve the following radical equations for x, and simplify: (5)√(x+1)^2 = 2

OpenStudy (anonymous):

if that's a fifth root, raise both sides to 5th power. (x+1)^2 = 32. +/- square root of 32 and simplify etc

OpenStudy (anonymous):

If not the 5th root: Divide both sides by 5 \[\sqrt{(x+1)^2} =2/5\] x+1 = 2/5 subtract 1 on both sides x = -3/5